OpenStudy (anonymous):

how do you solve x^4-625 = 0

OpenStudy (anonymous):

you can start by substituting \(x^2=u\)

OpenStudy (anonymous):

then solve for "u" (2 values) and then for "x" (2 values for each x, so four values in all)

OpenStudy (anonymous):

It can be written as: \[(x^2)^2 - (25)^2= = 0\] And you can use: \[x^2 - y^2 = (x-y)(x+y)\]

OpenStudy (anonymous):

(x-5)(x+5)

OpenStudy (anonymous):

Look it carefully and tell where it is x or x^2??

OpenStudy (anonymous):

try making it simpler \[u^2-625=0\\ \implies u=\pm\sqrt{625}=\pm25\\ x^2=25\qquad\qquad \boxed{x^2=-25}\;\text{(no real solutions, imaginary)}\\ \boxed{x = \pm 5} \]

OpenStudy (anonymous):

Then is it: (x^2+5)(x^2-5)

OpenStudy (anonymous):

No, that is 25^2 so, It would be 25.

OpenStudy (anonymous):

\[(x^2)^2 - (25)^2= 0 \implies (x^2 - 25)(x^2 + 25) = 0\]

OpenStudy (anonymous):

lets make it even simple \[zx^4-625=(x-25)(x+25)=(x-5)(x+5)(x^2+25)\]

OpenStudy (anonymous):

*x^2 @satellite73